OpenStudy (saadi101):

Solve using substitution: -3x + y = 1 and 4x + y = 8. Put your answer in coordinate form.

OpenStudy (jtug6):

Solve for one variable of either equation. How about the first? what does y equal for the first equation?

OpenStudy (saadi101):

I'm not sure how I go about doing any of this. Could you walk me through it I have 3 more questions like it

OpenStudy (jtug6):

Well solve for y for eqn 1. Add 3x to both sides perhaps? What does that give you for -3x + y = 1 if you add 3x to both sides?

OpenStudy (saadi101):

I don't understand explaining this in text, could you write out what your doing?

OpenStudy (jtug6):

\[-3x + y = 1\] and we're adding 3x to both sides

OpenStudy (jtug6):

\[-3x + 3x + y = 1 + 3x\] right?

OpenStudy (saadi101):

okay, yea

OpenStudy (jtug6):

so now you have \[y = 1 + 3x \] because -3x + 3x cancel right?

OpenStudy (saadi101):

right

OpenStudy (jtug6):

Now you know what y is, right? Well we have a 2nd equation which is: \[4x + y = 8\]

OpenStudy (jtug6):

So, if we know what y is what can we do to that equation?

OpenStudy (jtug6):

We're doing substitution, yes?

OpenStudy (saadi101):

right

OpenStudy (jtug6):

We know that y was equal to 1 + 3x so y has a new meaning too. It isn't just y anymore. We're substitution the what into y for \[4x + y = \]

OpenStudy (jtug6):

err \[4x + y = 8\]

OpenStudy (saadi101):

okay....

OpenStudy (jtug6):

Ok here i'll try to make it easier xD